OpenStudy (lena772):

Part A: Solve A = 9 over 2(x + 28) for x. (4 points) Part B: Determine the value of x when A = 135. (2 points) Part C: Solve –np – 40 > 10 for n. Show your work. (4 points)

OpenStudy (lena772):

Part A: A=\[\frac{ 9 }{ 2 }(x+28)\]

OpenStudy (shamil98):

9/2(x+28) = A multiply by the recprocal 2A/9 = x + 28 subtract 28 2A/9 - 28 = x

OpenStudy (shamil98):

B. Just evaluate it, A=135 put that in and solve for x. C. –np – 40 > 10 add 40 -np > 50 solving for n, divide by -p and flip the sign n < 50/-p

OpenStudy (lena772):

for b? (2*135)/9)-28=x (270/9)-28=x 30-28=x 2=x

OpenStudy (shamil98):

Yeah.

OpenStudy (anonymous):

@Lena772 i have the same question right now